# Unlocking the ‘Why’ of Buying: Introducing a New Dataset and Benchmark for Purchase Reason and Post-Purchase Experience

### v1

@@ 4.3 Characteristics of the Dataset
Table 5: The linguistic characteristics of our dataset. Higher Fleshch score indicates greater readability. Large lexical diversity score indicates higher diversity.
Word count Lex. diversity Fleshch
Product 184.84 0.68 49.71
User review 60.6 0.83 79.21
Explicit reason 11.39 0.97 64.63
Implict reason 10.16 0.96 48.75
Experience 22.19 0.93 51.16
5 A Benchmark of Recommendation Explaination
@@ 5.1 Task Definition
Given certain user information and the item information, a model is tasked to explain 1) why the user would purchase the item (i.e., purchase reason) and 2) what would the user’s experience be after using the item (i.e., post-purchase experience). Furthermore, we investigate the potential benefits of including auxiliary information, such as the user’s actual rating for the given item, to the explainability of the tasked model. We are interested in whether models can perform better with the predicted rating as a hint. Additionally, we explore the impact of adding an auxiliary task – rating prediction.
@@ 5.2 Experiment setup
We again choose large language models (LLMs) as the tasked models for benchmarking. It has been shown that explanations generated by LLMs are significantly favored by human raters over small models Liu et al. (2023). Specifically, we employ Gemini Ultra Gemini Team (2023) as the LLM for various tasks in a zero-shot setup (see Table 16 in Appendix for the basic prompt). We evaluate model generated explanations against the “ground-truth” explanations provided in our dataset. We use a variety of metrics to measure the generation performance: BLEU Papineni et al. (2002), Rouge1, Rouge2, and RougeLsum Lin (2004).
@@ 5.3 User and Item Representation
We are interested in learning the impact of different representation methods for the user and the item. To this end, we enhance our explanation dataset with past reviews for each user and each item. We base our analysis on our main task, task 1, defined in Section 5.1.
For user representation, we consider two options: 1) UserReview, composed of the user’s reviews written before purchasing the given item, and 2) UserProfile, where an LLM summarizes the same past reviews as the user profile (see Table 15 in Appendix for the prompt). For product representation, we explore three options: 1) Item, including the item’s title and description, 2) ItemReview, which additionally incorporates past reviews written by other users before the given user’s purchase, and 3) ItemProfile, where replaces the raw past reviews as an LLM generated summary. For past reviews, we select up to 10 of the most recent past reviews, with a maximum limit of 8k tokens.
As presented in Table 6, we see that representing users by their raw past reviews and items by their metadata is most effective in purchase reason generation. Post-purchase experience generation follows similar trend, but have a slight performance improvement by incorporating raw past reviews (Table 7). For users, summarizing their past reviews has the risk of losing personal specific information, while for items, including all reviews of other users may bring irrelevant information. Further research is needed to extract useful information from such noisy data. In the following experiments, we report the results using UserReview+Item combination.
Table 6: Performance on purchase reason generation using different user and item representations.
Method BLEU R1 R2 RLsum
UserReview-ItemReview 5.91 22.13 7.71 20.21
UserReview-Item 6.46 22.14 8.45 20.38
ItemReview 4.15 17.68 6.00 16.42
UserProfile-ItemProfile 4.64 17.95 5.71 16.66
UserReview-ItemProfile 3.88 15.44 5.05 14.48
UserProfile-ItemReview 5.25 8.67 2.94 7.99
UserProfile 5.24 18.71 6.51 17.32
Table 7: Performance on post-purchase experience generation using different user and item representations.
Method BLEU R1 R2 RLsum
UserReview-ItemReview 3.99 21.59 5.96 16.53
UserReview-Item 3.66 21.35 5.34 16.21
ItemReview 3.39 21.06 5.00 15.77
UserProfile-ItemProfile 3.70 21.25 5.18 15.94
UserReview-ItemProfile 3.51 20.35 4.60 15.17
UserProfile-ItemReview 3.89 9.36 2.54 7.2
UserProfile 3.85 20.78 4.92 15.94
@@ 5.4 Impact of task formulations
We study the impact of the auxiliary information and an auxiliary task. They respectively correspond to task 2 and task 3 defined in Section 5.1. The results, presented in Table 8, revealed no significant performance differences between the various task formulations. This finding indicates that LLMs exhibit robustness to variations in auxiliary information and tasks when operated in zero-shot settings.
Table 8: Performance of explanation generation based on different task formulations.
Task BLEU R1 R2 RLsum
@@
Task 1 6.46 22.14 8.45 20.38
Task 2 6.51 21.96 8.35 20.12
Task 3 6.44 21.77 8.34 19.97
@@
Task 1 3.66 21.35 5.34 16.21
Task 2 3.74 21.75 5.55 16.36
Task 3 4.49 22.41 6.00 17.62
@@ 5.5 Variation across Product Categories
We further break down the model performance by product category. As shown in Table 9 for purchase reason generation, the model excels in the “Electronic” and “Home and kitchen” categories, while it underperforms in “Clothing, Shoes & Jewelry” and “Books”. This aligns with our earlier findings in the auto-rater evaluation results (Section 4.2), where generated text for “Books” and “Clothing, Shoes & Jewelry” exhibited lower generality, as reflected in their "Non-personal" ratings. Conversely, the “Electronic” and “Home and kitchen” categories show higher generality. Consequently the explanation generation task is easier in these domains. Table 11 (in Appendix) shows the results for post-purchase experience generation with similar trend.
Table 9: Purchase reason generation performance per product category. Fashion denotes Clothing, Shoes and Jewelry.
Category BLEU R1 R2 RLsum
Overall 6.46 22.14 8.45 20.38
Books 5.36 19.44 6.68 17.50
Electronics 8.74 27.10 12.31 25.32
Home, Kitchen 7.21 24.17 9.35 22.54
Sports, Outdoors 6.15 22.04 8.02 20.27
Fashion 4.87 18.80 5.70 17.05
@@ 5.6 Discussions
Table 10 shows two example purchase reasons generated by our model. We observe that the model performance is contingent on the availability of purchase reason information within product metadata and the level of personalization in user reviews. When product metadata provides comprehensive insights into purchase reasons and user reviews are less individualized, the model exhibits strong predictive capabilities. However, the model’s performance degrades when attempting to generate purchase reasons that are more personalized, which may not be simply available from item metadata. This limitation calls for the development of models capable of constructing more comprehensive personal profiles to better predict such individualized purchase reasons.
@@
Example where the model generates a good explanation.
Product title: DEDC 1999-2007 Right Passengers Side Power Towing Mirrors Fit Ford Super Duty F250 F350 F450 1999 2000 2001 2002 2003 2004 2005 2006 2007
Ground-truth purchase reason: Needed a replacement right-side mirror for their Ford Super Duty truck (year 1999-2007)
Predicted purchase reason: User Anonymous may need replacement mirrors for their Ford Super Duty truck.
Example where the model generates a poor explanation.
Product title: The Whole Bible Story: Everything That Happens in the Bible in Plain English
Ground-truth purchase reason: To read the Bible in a condensed format, allowing the reader to easily navigate to specific chapters.
Predicted purchase reason: Interest in religion and biblical studies
@@ 6 Conclusion
We introduce a novel task of purchase reason explanation, aiming at better capturing what affects a user’s decision to purchase a product. We propose an LLM-based approach to generating a high quality, personalized dataset that consists of textual explanations of purchase reasons and post-purchase experiences based on user reviews. As the first of its kind, we demonstrate the dataset’s value by benchmarking it against purchase reason and post-purchase experience generation tasks. To empower further research, we release all scripts and prompts used for dataset creation and benchmarking.
With this new dataset, it is interesting to benchmark more explainable recommendation or explanation generation models, especially by refining user and item representations. Our results indicate that relying solely on all past reviews can introduce noise. It will be interesting to explore whether limiting the past reviews that are more similar to the item or the user could be beneficial, aligning with the ideas in retrieval augmented generation Gao et al. (2024). We leave this as future work.
@@ 7 Limitation
We propose a LLM-based solution for explanation dataset generation and auto evaluation. We validate this approach with Gemini Ultra and Amazon review dataset. Our approach is general, however, the performance with smaller models (e.g., Gemini Nano) and other model families (e.g., GPT-4, Claude, LLaMA) is still an open question. Similarly, we have not explored the generalization of our approach in reviews from other domains such Yelp and TripAdvisor.
@@ Appendix A Appendix
Table 11: Post-purchase experience generation performance per product category. Fashion denotes Clothing, Shoes and Jewelry.
Category BLEU R1 R2 RLsum
Overall 3.66 21.35 5.34 16.21
Books 3.05 19.71 3.94 14.85
Electronics 4.17 23.10 6.55 17.60
Home, Kitchen 4.16 22.52 6.22 17.29
Sports, Outdoors 3.79 21.89 5.73 16.51
Fashion 3.71 21.33 5.50 16.22
@@
You are a customer engagement specialist at Amazon, please analyze:
1. explicit_purchase_reason: why this customer purchased a product based on their reviews on Amazon. Describe in detail the thought processes before the purchase. Leave null if not mentioned.
2. implicit_purchase_reason: why this customer purchased this product, not mentioned in the reviews, can be from the product description.
3. purchase_reason_explanation: why do you think this is the purchase reason.
4. post_purchase_experience: how did the product meet this user’s expectation, describe it in 2 to 3 lines.
Be as specific and relating to personal context as much as possible.
Analyze purchase reason and experience based on this product review and product information.
@@
Please answer in json format, for example:
"explicit_purchase_reason": "…….."
"implicit_purchase_reason": "…….."
"purchase_reason_explanation": "…….."
"post_purchase_experience": "…….."
@@
You are a customer engagement specialist manager at Amazon and your specialist is trying to write a summary of why a specific purchase happened. Please assess the summary using the following criteria:
personal: Answer "Good", "Average" or "Bad", if the summary contains adequate personal context and is not too generic.
completeness: Answer "Yes" or "No" if most of the purchase reasons are covered in the summary answer "Yes".
completeness_reason: leave null if most of the reasons are covered.
personal: Answer "Good" or "Bad", if the answer contains adequate personal context and is not too generic.
personal_evidence: show evidence that this contains adequate personal context and is not too generic.
hallucination: Answer "Hallucination" or "Factual", are there any completely irrelevant information introduced, that are not implied in the product information or user review
hallucination_reason: if there are irrelevant information not implied in the product information or user review, show evidence
incorrect: Answer "Yes" if the summary contains information from after the purchase, else "No"
incorrect_reason: what information is from after the purchase?
Please answer in json.
Example assessment
@@
"incorrect": "Yes",
"incorrect_reason": "…"
Now let’s take a look at the following specialist’s task and summary and provide the assessment:
@@
Specialist summary of purchase reason:
@@
You are a customer engagement specialist manager at Amazon, your specialist is trying to write a summary of user post purchase sentiment.
Please assess the summary using the following criteria:
personal: Answer "Good" or "Bad", if the answer contains adequate personal context and is not too generic.
personal_evidence: show evidence that this contains adequate personal context and is not too generic.
hallucination: Answer "Hallucination" or "Factual", are there any completely irrelevant information introduced, that are not implied in the product information or user review
hallucination_reason: if there are irrelevant information not implied in the product information or user review, show evidence
completeness: Answer "Yes" if all the post purchase sentiments are covered in the summary and "No" otherwise.
completeness_reason: what are the missing sentiments?
Please answer in json.
Example assessment
"personal": "Good",
"personal_evidence": "…"
"hallucination": "Hallucination"
"hallucination_reason": "…"
"completeness": "Yes"
@@
Example assessment
"personal": "Average",
"personal_evidence": "…"
"hallucination": "Factual"
"hallucination_reason": "…"
"completeness": "No"
"completeness_reason": "…"
Now let’s take a look at the following specialist’s task and summary and provide the assessment:
@@
Specialist summary of post-purchase use experience:
Actual post-purchase use experience.
@@
Table 16: Prompt used to conduct task 1, purchase reason generation.
This is a user who left some past reviews on Amazon, please take a look at this user’s past review history on other products and predict this user’s purchase reason for this product and post purchase sentiment:

### v2

@@ 4.3 Characteristics and Differences to Other Datasets
Table 5: The linguistic characteristics of our dataset and two other explanation datasets. Larger type-to-token (TTT) score indicates higher lexical diversity.
Avg word count TTT
Product 156.33 0.058
User review 63.79 0.053
Purchase reason 11.39 0.140
Post-purchase Experience 22.19 0.048
EXTRA Li et al. (2021a) 4.96 0.006
P5 Geng et al. (2022) 25.33 0.014
4.4 Generalization of LLM Solution
To assess the broader applicability of our proposed LLM-based solution, we evaluate its performance using GPT-4 Turbo ((gpt-4-0125-preview) and Gemma-7B (1.1) on the same 100 review sample used for auto-rater evaluation. The quality of the extracted purchase reason and post-purchase experience data, as measured by the Gemini auto-rater (Table 6), reveals that larger models generally perform comparably, except Gemini (93.1%) outperforms GPT-4 (86.3%) in identifying complete purchase reasons. The smaller Gemma model is more prone to hallucination, introducing irrelevant information in 3.9% and 2.9% of purchase reason and post-purchase experience extractions, respectively. Additionally, Gemma’s post-purchase experience completion rate (88.2%) falls behind the other two models (94.1% for Gemini and 95.1% for GPT-4).
Table 6: The auto-rater results on datasets generated by different models. Hall./Compl./Type. denotes hallucination/completeness/type match, respectively.
Purchase Reason Experience
Model Hall. Compl. Hall. Compl. Type.
Gemini 1.0% 93.1% 0 94.1% 100%
GPT-4 2.0% 86.3% 0 95.1% 97.1%
Gemma 3.9% 87.3% 2.9% 88.2% 100%
5 A Benchmark of Recommendation Explanation
@@ 5.1 Task Definition
Given certain user information and the item information, a model is tasked to explain 1) why the user would purchase the item (i.e., purchase reason) and 2) what would the user’s experience be after using the item (i.e., post-purchase experience). Furthermore, we investigate the potential benefits of including auxiliary information, such as the user’s actual rating for the given item, to enhance the model’s explainability. We are also interested in whether models can perform better with the predicted rating as a hint. As such, we explore the impact of adding an auxiliary task – rating prediction.
@@
Table 7: Performance of Gemini Ultra using different user and item representations (%). B/R/RL/BERT denotes BLEU, ROUGE, ROUGE-Lsum, BERTScore respectively.
Purchase Reason Post-purchase Experience
Method B1 B2 R1 R2 RL BERT B1 B2 R1 R2 RL BERT
UserReview-ItemReview 17.7 8.8 22.8 8.9 21.1 25.5 21.8 8.7 24.2 5.4 20.0 31.3
UserReview-Item 18.1 9.4 23.4 9.8 21.7 26.1 23.0 8.7 24.3 5.4 20.2 29.3
ItemReview 13.0 6.5 20.8 7.7 19.4 23.2 22.2 9.2 23.9 5.9 19.7 30.5
UserProfile-ItemProfile 5.6 0.8 6.5 0.7 6.2 7.3 19.6 5.5 20.2 3.0 16.6 24.6
UserReview-ItemProfile 4.8 0.9 6.3 0.9 6.0 7.8 19.7 5.4 20.0 3.0 16.3 24.2
UserProfile-ItemReview 2.3 0.2 2.7 0.2 2.5 6.0 8.3 2.2 8.6 1.1 7.2 23.4
UserProfile-Item 7.0 1.6 8.8 1.6 8.3 10.2 20.3 5.4 20.1 3.0 17.0 23.9
@@ 5.2 Experiment setup
We again choose LLMs as the tasked models for benchmarking, given LLM-generated explanations are significantly favored by human raters over small pre-LLM models Liu et al. (2023). We evaluate model generated explanations against the “ground-truth” explanations provided in our dataset using various metrics. These include BLEU-1, BLEU-2 Papineni et al. (2002), ROUGE-1, ROUGE-2, and ROUGE-Lsum Lin (2004) to assess lexical similarity, and BERTScore Zhang et al. (2020) to measure semantic similarity. Our primary experimental LLM is Gemini Ultra, employed in a zero-shot setup (see basic prompt at Appendix Table 16). Additionally, we benchmark against GPT-4 Turbo, Gemma-7B, and a pre-LLM state-of-the-art model P5 Geng et al. (2022) for comparative analysis.
@@ 5.3 User and Item Representation
We are interested in learning the impact of different representation methods for the user and the item. To this end, we enhance our explanation dataset with past reviews for each user and each item. We base our analysis on task 1, our main task, as defined in Section 5.1.
For user representation, we consider two options: 1) UserReview, composed of the user’s reviews of other items written before purchasing the given item, and 2) UserProfile, where an LLM summarizes these reviews as the user profile (see Table 15 in Appendix for the prompt). For product representation, we explore three options: 1) Item, including the item’s title and description, 2) ItemReview, which additionally incorporates past reviews of this item written by other users before the given user’s purchase, and 3) ItemProfile, which replaces the past reviews with an LLM generated summary. For past reviews, we select up to 10 of the most recent past reviews, with a maximum limit of 8k tokens.
As shown in Table 7, representing users with their raw past reviews and items with their metadata is most effective for both generation tasks. Summarizing past user reviews risks losing personal details, while including reviews from other users may introduce irrelevant information. Further research is needed to extract useful information from such noisy data. In the following experiments, we report the results using UserReview+Item combination.
@@ 5.4 Impact of Task Formulations
We study the impact of an auxiliary information and an auxiliary task. They respectively correspond to task 2 and task 3 defined in Section 5.1. The results, presented in Table 8, reveal no significant performance differences between the various task formulations. This finding indicates that LLMs exhibit robustness to variations in auxiliary information and tasks when operated in zero-shot settings.
Table 8: Performance of Gemini Ultra based on different task formulations (%).
Task B1 B2 R1 R2 RL BERT
@@
Task 1 18.1 9.4 23.4 9.8 21.7 26.1
Task 2 18.1 9.4 23.4 9.8 21.6 25.8
Task 3 18.0 9.3 22.8 9.5 21.1 25.6
@@
Task 1 23.0 8.7 24.3 5.4 20.2 29.3
Task 2 22.9 8.9 24.4 5.5 20.1 29.7
Task 3 22.6 9.5 24.1 6.0 19.9 30.8
@@ 5.5 Variation across Product Categories
Additionally, model performance varies across product categories (Appendix Table 18). Post-purchase experience generation remains consistent, while purchase reason generation is stronger for “Electronics” and “Home and kitchen” compared to “Clothing, Shoes & Jewelry” and “Books”. This suggests that explanations are easier to generate for products with standardized purchase motivations, as opposed to those driven by subjective preferences and individual tastes.
5.6 Comparison Among Different Models
Table 9: Performance of all methods on 1k sampled dataset (%). 1s denotes for one-shot. Gemini Ultra is statistically significant better than other methods via a two-sided paired t-test ($p<0.01$) on all metrics.
B1 B2 R1 R2 RL BERT
Purchase reason
Gemini 16.5 8.4 21.8 9.2 20.1 25.2
GPT-4 (1s) 14.3 5.8 18.0 5.4 16.1 19.2
Gemma (1s) 8.2 0.1 6.6 0.1 7.9 10.6
P5 1.2 0.06 2.7 0.06 2.6 8.5
Post-purchase experience
Gemini 22.9 8.9 24.1 5.7 20.0 28.5
GPT-4 (1s) 20.1 5.6 21.4 3.1 16.8 23.5
Gemma (1s) 19.6 2.5 14.5 1.6 16.6 21.8
P5 0.6 0.05 4.9 0.18 4.7 7.1
To further evaluate our dataset, we additionally benchmark it using GPT-4-Turbo, Gemma-7B and a pre-LLM state-of-the-art model P5 on a randomly sampled 1,000 reviews (see results in Table 9). P5, a T5-based model trained on five recommendation tasks (e.g., rating prediction, explanation generation) using Amazon reviews, frequently generates short, generic explanations (e.g., “great product”). This aligns with prior findings Liu et al. (2023) and results in poor performance regardless of whether the explanations are classified as purchase reasons or post-purchase experiences.
Table 10: Performance on 1k sampled dataset with ground truth generated by GPT-4 (%). Gemini Ultra is statistically significant better than GPT-4 via a two-sided paired t-test ($p<0.01$) on all metrics.
B1 B2 R1 R2 RL BERT
Purchase reason
Gemini 15.8 6.7 21.7 6.7 19.7 21.9
GPT-4 (1s) 14.6 5.2 19.6 5.0 17.4 17.5
Post-purchase experience
Gemini 24.5 9.3 24.5 5.1 19.9 26.8
GPT-4 (1s) 23.4 6.8 23.4 3.4 17.8 22.8
Initially, GPT-4 performed significantly worse than Gemini with the same zero-shot prompt. Adding a one-shot example improves GPT-4’s performance, but it still performs worse than Gemini. This raises concerns about potential bias introduced by using an LLM generator (Gemini Ultra) for the ground truth. To address this, we use GPT-4 to generate the ground truth purchase reasons and post-purchase experiences and re-evaluate performance (Table 10). The results consistently demonstrate Gemini’s superior performance, regardless of the ground truth source. Finally, we observe that the smaller Gemma-7B model, even with a one-shot prompt, lags significantly behind the larger models. Future research could explore fine-tuning smaller models to bridge this performance gap.
@@ 5.7 Discussions
We observe that the model performance is contingent on the presence of relevant information within user’s previous reviews. Table 11 shows two typical examples of good and poor purchase reason generation by Gemini Ultra. In Example 1, where the user bought a navigation receiver for a Chevrolet Silverado truck, the model accurately deduces the purchase reason for the current product. However, in Example 2, the lack of relevant information in the user history forces the model to rely on product metadata, resulting in a generic purchase reason. This decline in performance is especially noticeable when the actual purchase reason is highly personalized or differs significantly from the product metadata.
This limitation calls for the development of more capable models for better personalized purchase reason prediction. It’s crucial to develop models that can effectively identify relevant information from user’s past reviews. Furthermore, since our current setup utilizes only the 10 most recent reviews to represent users, it’s worth investigating whether incorporating a larger portion of relevant user history could enhance model performance. This can be achieved through models with extended context capabilities (e.g. Gemini 1.5 Pro Gemini Team (2024), or through retrieval-augmented generation techniques Gao et al. (2024). We hope the dataset developed by this paper can inspire more research towards these directions.
@@
Example 1 (the model generates a good explanation) Product title: Pedal Commander throttle response controller PC65 for GM models 2007 and newer - get increased performance or save fuel up to 20% - Available for Silverado, Sierra, Tahoe, Escalade, Yukon, Suburban Ground-truth purchase reason: To increase the performance of their 2016 Chevy Silverado. Predicted purchase reason: To improve the performance of their Chevrolet Silverado truck.
Model’s rationale: The user’s review of the Alpine Electronics X110-SLV indicates that they value sound quality and performance in their vehicle systems.
Example 2 (the model generates a poor explanation) Product title: QALO Men’s Charcoal Step Edge Q2X Silicone Ring Size 08 Ground-truth purchase reason: Look for a comfortable ring for work and yard work. Predicted purchase reason: User Anonymous may be looking for a durable and stylish accessory.
Model’s rationale: The user has not left any reviews on similar products, but the product description highlights its durability and style.
@@ 6 Conclusion
We introduce a novel task of purchase reason explanation, aiming at better capturing what affects a user’s decision to purchase a product. We propose an LLM-based approach to generating a high quality, personalized dataset that consists of textual explanations of purchase reasons and post-purchase experiences based on user reviews. Our approach demonstrates strong generalization ability, yielding strong performance across diverse large language models, including Gemini Ultra, GPT-4 Turbo, and Gemma-7B. As the first of its kind, we showcase the dataset’s value by benchmarking it against purchase reason and post-purchase experience generation tasks. To empower further research, we release all scripts and prompts used for dataset creation and benchmarking.
This new dataset opens up exciting possibilities for building better explanation generation models, particularly by refining user and item representations. This includes exploring the direct incorporation of more past reviews through models with extended context windows or leveraging retrieval-augmented generation techniques to extract relevant information. Additionally, our study reveals a significant performance gap between smaller and larger models in a few-shot setting. A natural extension is to explore if fine-tuning a smaller model can effectively close this gap. We leave these as future work.
@@ 7 Limitation
We propose a LLM-based solution for explanation dataset generation and auto evaluation. We validate this approach using few-shot learning with three LLMs (Gemini Ultra, GPT-4 Turbo, and Gemma-7B) on Amazon reviews. While our approach is general, performance with other model families (e.g., Claude, LLaMA) and smaller fine-tuned models remains an open question. Similarly, we have not explored the generalization of our approach in reviews from other domains such as Yelp and TripAdvisor.
Our benchmarking evaluation utilizes lexical and semantic similarity metrics (BLEU, ROUGE, BERTScore) to measure similarity between model-generated explanations and the ground truth. While informative, these metrics have limitations and do not fully capture the nuances of LLM-generated text. An automated LLM evaluation system could potentially address these limitations. Additionally, a user study or online A/B test would be valuable for assessing the real-world effectiveness of purchase reason explanations.
@@ Appendix A Appendix
A.1 Prompts
This section shows a few prompts used by our LLMs, including
  • The prompt used for extracting explanations from user reviews (Table 12).
  • The prompt used by LLM auto-rater to evaluate extracted purchase reason (Table 13) and post-purchase experience (Table 14).
  • The prompt to summarize user past reviews (Table 15).
  • The prompts to generate explanations in both zero-shot (Table 16) and one-shot (Table 17) setups.
@@ A.1 Prompts
You are a customer engagement specialist at Amazon. Given a product and a customer review, please identify the customer’s purchase reason and post-purchase experience:
1. explicit_purchase_reason:The reason for purchasing the product that is directly mentioned in the review. Describe in detail the thought processes before the purchase. Leave null if not mentioned.
2. implicit_purchase_reason: The reason for purchasing the product that is not explicitly mentioned in the review but can be inferred from the review and the product description. Leave null if not mentioned.
3. purchase_reason_explanation: Briefly justify your reasoning behind the identified explicit and implicit purchase reasons.
4. post_purchase_experience: Summarize in 2-3 lines how the product met (or didn’t meet) the customer’s expectations, based on the review.
Please be as specific and relating to the customer’s personal context as much as possible.
Please answer in valid json format, for example:
"explicit_purchase_reason": "……..",
"implicit_purchase_reason": "……..",
"purchase_reason_explanation": "……..",
"post_purchase_experience": "…….."
Here are two examples: Product:
Suncast PB6700 Patio Bench, Light Taupe
Suncasts 50 gallon patio bench provides comfortable seating along with convenient storage. It is perfect for storing gardening supplies, patio accessories and more. This patio bench is decorative and functional and will look great in your backyard.
Customer Review:
Good but….
I bought this to use as a toy box and seating for my kids. After putting it together which is more tricky than you first think I put toys in it and let my kids play with it then I noticed on the instructions it says not for kids or to be used as a toy box but I’m not sure why. It holds all my kids toys and they love it and as long as I don’t let them stand on it or jump on it I think it’s perfect. It’s not as sturdy as I thought but let’s hope it holds up and lasts a while
Expected Answer:
"explicit_purchase_reason": "To store kids’ toys and provide seating for them.",
"implicit_purchase_reason": "50-gallon storage capacity",
"purchase_reason_explanation": "The customer needed a solution that could both store toys and serve as a seating area for their children.",
"post_purchase_experience": "The bench met the customer’s expectations for storage and seating. However, it was not as sturdy as they had anticipated."
Product:
The World of Beads
Muto is a Kodansha International author.
Customer Review:
Five Stars
AWESOME AWESOME BOOK..:)
Expected Answer:
"explicit_purchase_reason": null,
"implicit_purchase_reason": null,
"purchase_reason_explanation": "The review does not provide any information about the customer’s purchase reasons.",
"post_purchase_experience": "The customer was very satisfied with the book, describing it as ’AWESOME’ and ’Five Stars’."
Please analyze purchase reason and post-purchase experience based on this product review and product information.
@@
As a customer engagement team leader at Amazon, your task involves evaluating a summary written by a specialist about why a certain purchase was made.
You will analyze the summary based on the provided product information and customer review, using these criteria:
1. completeness: Answer "Yes" or "No". "Yes" if the summary successfully captures the majority of reasons behind the purchase; otherwise, "No".
2. completeness_reason: Provide a concise explanation for your assessment of the summary’s completeness.
3. hallucination: Answer "Hallucination" or "Factual". "Hallucination" if the summary includes any unrelated details not mentioned or implied by the product information or the customer review; otherwise, "Factual".
4. hallucination_reason: Provide a concise explanation for your assessment of the summary’s hallucination.
5. correctness: Answer "Yes" or "No". "Yes" if the summary exclusively focuses on pre-purchase information without discussing any post-purchase experiences; otherwise, "No".
6. correctness_reason: Provide a concise explanation for your assessment of the summary’s correctness.
Please respond using a valid json format, for example:
@@
"correctness": "Yes",
"correctness_reason": "…"
Now, please evaluate the following summary based on the above criteria:
@@
Specialist’s summary of the reasons for purchase:
@@
As a customer engagement team leader at Amazon, your task involves evaluating a summary written by a specialist about a user’s experience after purchasing a product.
You will analyze the summary based on the provided customer review and product information, using these criteria:
1. completeness: Answer "Yes" or "No". "Yes" if the summary successfully captures the majority of the customer’s post-purchase experience; otherwise, "No".
2. completeness_reason: Provide a concise explanation for your assessment of the summary’s completeness.
3. hallucination: Answer "Hallucination" or "Factual". "Hallucination" if the summary includes any unrelated details not mentioned or implied by the product information or the customer review; otherwise, "Factual".
4. hallucination_reason: Provide a concise explanation for your assessment of the summary’s hallucination.
Please respond using a valid json format, for example:
"hallucination": "Hallucination",
"hallucination_reason": "…",
"completeness": "Yes",
@@
Now, please evaluate the following summary based on the above criteria:
@@
Specialist’s summary of post-purchase experience:
Actual post-purchase experience.
@@
Table 16: Zero-shot prompt used to conduct task 1.
This is user Anonymous who left some past reviews on Amazon. Please take a look at this user Anonymous’ past review history on other products and predict user Anonymous purchase reason for this product and post purchase sentiment